# Diffusing in Someone Else’s Shoes: Robotic Perspective Taking with Diffusion

### v1

@@ Abstract
Humanoid robots can benefit from their similarity to the human shape by learning from humans. When humans teach other humans how to perform actions, they often demonstrate the actions and the learning human can try to imitate the demonstration. Being able to mentally transfer from a demonstration seen from a third-person perspective to how it should look from a first-person perspective is fundamental for this ability in humans. As this is a challenging task, it is often simplified for robots by creating a demonstration in the first-person perspective. Creating these demonstrations requires more effort but allows for an easier imitation. We introduce a novel diffusion model aimed at enabling the robot to directly learn from the third-person demonstrations. Our model is capable of learning and generating the first-person perspective from the third-person perspective by translating the size and rotations of objects and the environment between two perspectives. This allows us to utilise the benefits of easy-to-produce third-person demonstrations and easy-to-imitate first-person demonstrations. The model can either represent the first-person perspective in an RGB image or calculate the joint values. Our approach significantly outperforms other image-to-image models in this task.
@@ I Introduction
Imitation Learning or Learning from Demonstrations (LfD) is a learning mechanism often encountered in humans because of how direct a demonstration can transmit information. If an action is taught by explaining it, a further level of abstraction is needed, which direct demonstration can bypass. However, it does require other abilities that can be found in humans but are more complex than they might seem. When learning from a demonstration, correctly perceiving the demonstration is the first step. Then, this perceived action has to be transferred to ourselves so that we can imitate it. One of the necessary abilities for this process is perspective-taking.
Visual and spatial perspective-taking is the ability to see someone and imagine what they are seeing. This ability requires an understanding of the objects that are in someone’s field of view and understanding that they will see the objects differently. The objects will have to be transferred from the position, shape, and size one person is seeing them to the position, shape, and size another person is seeing them. In humans this ability develops between the ages 3-4 [1]. In this paper, we show that diffusion models can also learn perspective-taking. This capability allows our approach to directly find the joint configurations necessary to imitate behaviour as seen from a third-person view. It combines the need to perceive the demonstration and to transfer it to our perspective and even body when predicting joint values.
Figure 1: On the left, the figure shows a demonstration in the simulation and the resulting imitation on the real robot on the right.
@@
Apart from directly learning from imitation, our approach can also transfer images from the third-person view to the first-person view. This functionality can be used to generate data for the many existing imitation learning approaches that already exist. ”Unfortunately, hitherto imitation learning methods tend to require that demonstrations are supplied in the first-person […] While powerful, this kind of imitation learning is limited by the relatively hard problem of collecting first-person demonstrations” Stadie et al. [2]. Our approach is able to generate the first-person view from the third-person view, thereby removing the challenge of collecting first-person view data. We also attempt to directly imitate the demonstrations from the third-person view.
@@
The field of image generation has been quickly improving over the past years through GANs and diffusion models [3, 4, 5, 6]. To generate the first-person view, we make use of the diffusion models in particular. Diffusion models work by purposely putting noise on a desired output and then iterative denoising it again. Through this process, the model learns to generate the desired output from pure noise by detecting general patterns in the desired outputs. The model also learns to build upon its last prediction as the denoising process is iterative, so it also develops the ability to perceive the given images quite well. To direct the generated output, a condition can be added to the diffusion model. This can be a label or a prompt. However, diffusion has also been used for tasks in which the condition is what would be the input in classical models, such as an image for object detection [7]. We follow this trend of having a high-dimensional and complex condition, that should result in a specific generated image. The condition displaying the third-person image is given to the model in various stages to allow it to generate a fitting first-person image.
Our main contributions are:
@@
  • A novel architecture to generate images in first-person perspective from images in third-person perspective for a humanoid robot.  [^1] We plan to publish our code at https://www.inf.uni-hamburg.de/en/inst/ab/wtm/research/software.html
  • Our model outperforms competing approaches such as pix2pix and CycleGAN.
  • The approach can directly infer and translate the first-person pose from the third-person image as shown in Fig. 1., through small modifications to the architecture.
@@
  • We publish a new dataset with pairs of images for third-person and first-person perspectives. [^2] We plan to publish our dataset at https://www.inf.uni-hamburg.de/en/inst/ab/wtm/research/corpora.html
@@ II Related Work
Isola et al. [8] demonstrated that generative image-to-image models can solve a multitude of tasks. Ranging from changing an image from black and white to coloured, to generating an image just from the edges or turning segmentation labels into a scene image. They achieved this using conditional Generative Adversarial Networks (GANs). Further advancements in image-to-image models have been made by incorporating special losses such as cycle consistency loss, which ensures specific outputs for given inputs [9]. Going from the third-person perspective to the first-person perspective can also be done using their model, the results tend to be not as good as for some of the other tasks.
@@
Specialised models have been developed for this specific task of perspective-taking. To focus the model on a general understanding of the demonstration, the kernel sizes of the convolutional layers were increased, and self-attentive layers were added to help with understanding the relations between different parts of the input images. As a demonstration is often a continuous video, past frames can be used to improve the performance [10]. In a more general approach towards perspective transfer without the direct relation to robots and imitation learning, Liu et al. proposed the Parallel GAN (P-GAN) architecture [11]. By doing both a transfer from first-person to third-person and, at the same time, a transfer from third-person to first-person, they were able to outperform other models in this task, such as X-Seq or X-Fork [12] where the GANs were not working in parallel.
@@
In base diffusion models, a neural network architecture is used that starts with an image input, encodes it and decodes it again to reconstruct it. The key is that different levels of noise are put on the input image. The model, therefore, has to learn to either predict which parts of the image are noise, or what it should look like without the noise [3]. Once trained, the model can iteratively generate images by progressively removing noise levels encountered during training. As diffusion models possess a similar purpose to GANs in generating images, it does not come as a surprise that they can be used for many similar tasks, including image-to-image translations. Pix2Pix-zero is one such model, which is able to do zero-shot image translation by changing an image along an ”edit direction” [9]. Other diffusion approaches focus more on generating images from text prompts, layouts or semantic synthesis [13].
@@
Similar to GANs, diffusion models can also be conditioned to generate specified output. In diffusion models, there is always one natural way of condition, which is the time step. This is a parameter that tells the network what level of noise it can expect so that it can adjust to that. Additional conditions, such as a text embedding to direct the output, can be added to this time step [14]. The conditioning has also been used for tasks a bit further away from generating images, such as action segmentation [15] or object detection [7]. Here, the output is the bounding box or frame label for the condition input. These works use an encoder-decoder structure instead of a U-Net, and the condition is added to the input of the decoder layers.
When it comes to imitation, movements play a big role. For robots, Inverse Kinematics solvers tackle the task of moving kinematic body parts to predefined targets, a problem that can quickly increase in complexity with the Degrees of Freedom (DoF) the body part possesses [16]. There can be infinite solutions for one inverse kinematic problem, making inverse kinematics very challenging [17]. Deep learning methods, including generative models, have also been adapted to this task, generating the movement necessary to reach the defined position. Similar to image-to-image tasks, the relation between the in- and output is crucial, making losses like the cycle consistency loss useful for this task [18]. Some approaches have also experimented with further changes to the GAN architecture, changing the input, the costs, the output or ensembling multiple networks [19].
@@
Learning from demonstration is utilised to get past many challenges of machine learning. From a demonstration, much information about how to solve a task can be gained, and the demonstrated solution can be used to recreate that success. Imitation learning can be used to allow robots to learn movement planning. As going from RBG images to joints can be difficult, different kinds of observing the demonstrations have been used, such as voxels [20] or point clouds [21]. Although, for human data, pose detectors exist that are able to find key points of the poses directly from RGB images [22], demonstrations are often performed from a first-person perspective so that trajectories stay similar and the demonstration is closer to the imitation. To directly learn from a third-person perspective, the trajectories have to be translated to the first-person perspective, which can be achieved with the use of generative models [2]. Even just mirroring a pose seen in a mirror can be very challenging for a robot. Hart and Scassellati developed an approach that makes use of six models, an end-effector model, a perceptual model, a perspective-taking model, a structural model, an appearance model and a functional model for this challenge. Their loss was six times higher for demonstrations from third-person view compared to demonstrations from first-person view.[23].
@@ III Methods
Our approach is based on a conditioned diffusion model. In contrast to the model from Ho et al. [3], we use a full image as our condition, akin to how DiffusionDet [7] or Diffusion Action Segmentation [15] made use of diffusion models. The condition is an image captured from a camera opposite the robot, representing how a person sitting across from it would see the robot. The model then generates the scene as seen by the robot. As the approach uses diffusion, we also have another input, which during inference is simply noise and during training consists of the desired output with noise applied. The level of noise depends on a randomly chosen time step between 0 and 1000 for each sample. A cosine noise schedule calculates the amount of noise depending on the chosen time step [24]. The cosine schedule allows for a smoother transition from no noise to full noise compared to a linear noise schedule. While the network should mostly rely on the condition, the smoother transition can increase the benefits of the iterative denoising process.
Figure 2: Our architecture has two inputs: the condition, which is the third-person view of the robot in the top middle and the noised output during training or, as seen in the top left image, simply noise during inference. Then, the top left image shows the noise that is the input during inference, and the top right image shows the output of our model, the first-person view from the robot. Both input images are decoded, each through one of the downward paths. The decoded version of the noise is directly led into the upward path that decodes the image to the output, while the encoded version of the condition is led into the upward path through latent connections.
@@
The architecture of our model is shown in Fig. 2. The structure of the model largely follows a U-Net [25], however, there are no latent connections between the primary downward and upward leading paths of the U-Net. Instead, we added a secondary downward leading path that has the purpose of encoding our condition. This path is only connected through the latent connections to the rest of the architecture. Our model directly predicts the desired output, instead of trying to predict the noise, as proposed by Song et al. [4]. This decision was made as our goal lies more in correct predictions and accuracy than in the model’s ability to generate new ideas or varying outputs, as could be desired when generating images from prompts. During inference, we start with using randomly generated noise for our input and the condition, which is the RBG image taken from a different perspective providing the information needed to replicate the pose seen in it. Predicting the desired result allows us to utilise only 50 denoising steps, significantly accelerating the inference. Our model leverages both the information provided by the condition and that acquired during later steps of the iterative denoising process to generate its output.
@@
Both paths follow the same structure of convolution layers, attention layers and maxpooling layers. The first path ends in two convolution layers that do not exist in the second downward path. The upward path consists of convolutional layers, self-attention layers and upsampling layers in order to get back to the dimensions of the original input. The latent connections from the second downward path to the upward path always go from the self-attention layer to the upsampling layers. The time step is encoded and given in the linear layers of the ”Down” and ”Up” modules.
We can modify our architecture by adding one linear layer in order to encode the input from the form of the joint values, which are two arrays of length 13, to the same dimensions as our conditioning image, which is an RBG image with a resolution of $64\times 6464 × 64$ pixels. Additionally, one linear decoding layer is appended at the end, to convert the dimensions from $64\times 64\times 364 × 64 × 3$ into $2\times 132 × 13$ joint values. This allows us to generate joint states instead of an RGB image.
To create our dataset, we made use of the NICOL simulation introduced by Kerzel et al. [26]. We use two cameras in the simulation, one situated in the head of the robot and the second one positioned in front of it at a distance that allows us to fully capture the robot. The robot’s head is tilted downwards to ensure the actions are captured in the first-person view. The NICOL robot is a semi-humanoid robot without legs integrated into a collaborative workspace. We make the robot randomly assume poses for both of its arms. In each of the kinematic chains representing the arms, we have 13 joints. Five of these joints are in the hands, while the other eight are in the arms. The joint values range from minus pi to pi, which we scale to be between 0 and 1 for the training. After the poses have been taken, we take one image from both cameras and record the joint values from the robot. The end-effector position and rotation are also recorded in seven values for each of the hands, with the first three values representing the x, y, and z coordinates and the other four values representing the orientation. During the data collection, 10000 samples were recorded. For training, 80% of the dataset is used, while 20% is used for validation. All the images are RBG and have a height and a width of 64 pixels. The random poses come from a uniform distribution but are limited to stay in the visual range and within the ranges of the joints. There are no images where the arms are behind the robot’s back or otherwise obscured. The finger joints are always kept in the same position, so only eight of the thirteen recorded joint values per arm are changed. We have tried out multiple preprocessing steps, such as rotating, flipping, normalising and masking parts of the image.
@@
The network was trained for 100 epochs; as mentioned earlier, the noise applied to each training sample is chosen randomly, and there are 1000 different levels of noise possible. This extends the variance of our dataset, as the model is supposed to predict the same result regardless of the level of noise that was applied to the input. We utilised the Adam optimiser and the MSE loss function.
@@
In this end-to-end approach, the network manages to identify the three-dimensional pose of the robot from a two-dimensional RBG image and to transpose that pose from a third-person perspective to an egocentric perspective, which can be expressed through a generated RGB image or joint states.

### v2

# Diffusing in Someone Else’s Shoes: Robotic Perspective-Taking with Diffusion

@@ Abstract
Humanoid robots can benefit from their similarity to the human shape by learning from humans. When humans teach other humans how to perform actions, they often demonstrate the actions, and the learning human imitates the demonstration to get an idea of how to perform the action. Being able to mentally transfer from a demonstration seen from a third-person perspective to how it should look from a first-person perspective is fundamental for this ability in humans. As this is a challenging task, it is often simplified for robots by creating demonstrations from the first-person perspective. Creating these demonstrations allows for an easier imitation but requires more effort. Therefore, we introduce a novel diffusion model that enables the robot to learn from the third-person demonstrations directly by learning to generate the first-person perspective from the third-person perspective. The model translates the size and rotations of objects and the environment between the two perspectives. This allows us to utilise the benefits of easy-to-produce third-person demonstrations and easy-to-imitate first-person demonstrations. Our approach significantly outperforms other image-to-image models in this task.
@@ I Introduction
Imitation Learning or Learning from Demonstrations (LfD) is a learning paradigm often encountered in humans which efficiently transmits knowledge about processes. If alternatively an action is taught verbally by explaining it, a further level of abstraction is needed, which demonstrations can bypass. However, learning from demonstrations necessitates a set of sophisticated abilities. In humans, these abilities develops between the ages three and four [2]. When learning from a demonstration, correctly perceiving the demonstration is the first step. Then, this perceived action has to be transferred to our own perspective so that we can imitate it [1]. One of the necessary abilities for this process is perspective-taking.
Visual and spatial perspective-taking is the ability to see someone and imagine what they are seeing. This ability requires an understanding of the objects that are in someone’s field of view and understanding that they will see the objects differently. The objects will have to be transferred from the position, shape, and size as seen by one person to the position, shape, and size as seen by another person. In this paper, we looked into how a machine learning model can achieve similar abilities. Towards this purpose we developed a diffusion model that utilises images recorded from a third-person perspective to generate corresponding images from a first-person perspective, as shown in Fig. 1.
Figure 1: The results of our model for one example of each dataset. Top row is our dataset created with the simulated NICOL robot. An example of the EMIL dataset [3] which was recorded in a real-world scenario is in the second row. The third row is for the RH20T dataset [4] which was recorded in a more cluttered real-world scenario. The resolution is in accordance with the model’s input and output.
@@
The generation of first-person perspectives can be used to generate data for the many existing imitation learning approaches, that learn from first-person perspective demonstrations. This will remove a challenge described by Stadie et al.: “Unfortunately, hitherto imitation learning methods tend to require that demonstrations are supplied in the first-person […] While powerful, this kind of imitation learning is limited by the relatively hard problem of collecting first-person demonstrations.”
@@
The field of image generation has been quickly improving over the past years through Generative Adversarial Networks (GANs) [5] and diffusion models [6, 7, 8]. To generate the first-person perspective, we make use of a diffusion model. Diffusion models work by purposely putting noise on a desired output and then iteratively denoising it again. Through this process, the model learns to generate the desired output from pure noise by detecting and refining general patterns in the desired outputs. The model also learns to build upon its last prediction as the denoising process is iterative, so it also develops the ability to perceive the given images quite well. To direct the generated output, a condition can be added to the diffusion model, such as a label or a prompt. However, diffusion has also been used for tasks in which the condition is what would be the input in classical models, such as an image for object detection [9]. We similarly have a high-dimensional and complex condition, that should result in a specific generated image. The condition displaying the third-person image is given to the model in various stages to allow it to generate a fitting first-person image. Our main contributions are:
@@
  • We present a novel architecture to generate images in first-person perspective from images in third-person perspective for a humanoid robot, as shown in Fig. 2.
@@
  • We publish our code and a newly generated dataset with pairs of images for third-person and first-person perspectives. [^1] We plan to publish our code and dataset at https://github.com/knowledgetechnologyuhh /robotic-perspective-taking-with-diffusion
  • We evaluate our architecture on three different datasets.
  • Our model outperforms competing approaches such as pix2pix [10] and CycleGAN [11].
@@ II Related Work
Generative image-to-image models. Isola et al. [10] demonstrate that generative image-to-image models can solve a multitude of tasks. Ranging from changing an image from black and white to coloured, to generating an image just from the edges or turning segmentation labels into a scene image. They achieved this using conditional GANs. Further advancements in image-to-image models have been made by incorporating special losses such as cycle consistency loss, which ensures specific outputs for given inputs [11]. Going from the third-person perspective to the first-person perspective can also be done using their model, but the results tend to be not as good as for some of the other tasks, as seen in the results section.
@@
Perspective-Taking. Specialised models have been developed for this specific task of perspective-taking. To focus the model on a general understanding of the demonstration, the kernel sizes of the convolutional layers were increased, and self-attentive layers were added to improve the understanding of the relations between different parts of the input images. As a demonstration is often a continuous video, past frames can be used to improve the performance [12]. In a more general approach towards perspective-taking without the direct relation to robots and imitation learning, Liu et al. proposed the Parallel GAN (P-GAN) architecture [13]. By doing both a transfer from first-person to third-person and, at the same time, a transfer from third-person to first-person, they were able to outperform other models in this task, such as X-Seq or X-Fork [14] where the GANs were not working in parallel.
@@
Diffusion Models. In base diffusion models, a neural network architecture is used that starts with an image input, encodes it and decodes it again to reconstruct it. The key is that different levels of noise are put on the input image. The model, therefore, has to learn to either predict which parts of the image are noise, or what it should look like without the noise [6]. Once trained, the model can iteratively generate images by progressively removing noise levels encountered during training. As diffusion models possess a purpose similiar to GANs in generating images, it does not come as a surprise that they can be used for many similar tasks, including image-to-image translations. Pix2Pix-zero is one such model, which is able to do zero-shot image translation by changing an image along an “edit direction” [11]. Other diffusion approaches focus more on generating images from text prompts, layouts or semantic synthesis [15].
@@
Conditional Diffusion Models. Similar to GANs, diffusion models can also be conditioned to generate specified output. In diffusion models, the time step is inherently a part of the conditioning process. This parameter specifies the applied level of noise, allowing the network to correctly interpret the input in relation to this noise level. Additional conditions, such as a text embedding to direct the output, can be added to this time step [16]. The conditioning has also been applied to other tasks apart from image generation, such as action segmentation [17] or object detection [9]. In such cases, the output is the bounding box or frame label for the condition input. These works utilise an encoder-decoder structure, in which the condition is added to the input of the decoder layers. In contrast, U-Nets [18], which have additional skip connections, are commonly used for image generation tasks [16] because their skip connections can retain more spatial information.
Figure 2: Our architecture has two inputs: the condition, which is the third-person perspective of the robot in the top middle and the noised output during training or, as seen in the top left image, simply noise during inference. Then, the top left image shows the noise that is the input during inference, and the top right image shows the output of our model, the first-person perspective from the robot. Both input images are encoded, each through one of the downward paths. The encoded version of the noise is directly led into the upward path that decodes the image to the output, while the encoded version of the condition is led into the upward path through latent connections.
@@
Robotic Imitation. Learning from demonstration is utilised to get past many challenges of machine learning. From a demonstration, much information about how to solve a task can be gained, and the demonstrated solution can be used to recreate that success. Imitation learning can be used to allow robots to learn movement planning. As going from RBG images to joints can be difficult, different methods to record the demonstrations have been used, such as voxels [19] or point clouds [20]. Although, for human data, pose detectors exist that are able to find key points of the poses directly from RGB images [21], demonstrations are often performed from a first-person perspective so that trajectories stay similar and the demonstration is closer to the imitation. To directly learn from a third-person perspective, the trajectories have to be translated to the first-person perspective, which can be achieved with the use of generative models [22]. Even just mirroring a pose seen in a mirror can be very challenging for a robot. Hart and Scassellati [23] develop an approach that makes use of six models: an end-effector model, a perceptual model, a perspective-taking model, a structural model, an appearance model and a functional model for this challenge. Their loss was six times higher for demonstrations from a third-person perspective compared to demonstrations from a first-person perspective.
@@ III Methods
Figure 3: Qualitative comparisons between our model and pix2pix as well as CycleGAN. Five examples (A, B, C, D, E) are shown to illustrate the differences between the models. CycleGan has almost no variance between its output, pix2pix mostly seems to have the correct poses except for example C, but is often unable to correctly recreate the hands and finger structure. Our model has always the correct pose and fully recreates the hands. The resolution is in accordance with the model’s input and output.
III-A Architecture
@@
Our approach is based on a conditioned diffusion model. In contrast to the model from Ho et al. [6], we use a full image as our condition, akin to how DiffusionDet [9] or Diffusion Action Segmentation [17] made use of diffusion models for computer vision tasks. The condition is an image captured from a camera opposite the robot, representing how a person sitting across from it would see the robot. The model then generates the scene as seen by the robot. As the approach uses diffusion, we also have another input, which during inference is simply noise and during training consists of the desired output with noise applied. The level of noise depends on a randomly chosen time step between 0 and 1000 for each sample. A cosine noise schedule calculates the amount of noise depending on the chosen time step [24]. The cosine schedule allows for a smoother transition from no noise to full noise compared to a linear noise schedule. While the network should mostly rely on the condition, the smoother transition can increase the benefits of the iterative denoising process.
@@
The architecture of our model is shown in Fig. 2. The structure of the model largely follows a U-Net [18], however, there are no latent connections between the primary downward and upward leading paths of the U-Net. Instead, we added a secondary downward leading path that encodes our condition. This path is only connected through the latent connections to the rest of the architecture. Our model directly predicts the desired output, instead of trying to predict the noise, as proposed by Song et al. [7]. This decision was made as our goal lies more in correct predictions and accuracy than in the model’s ability to generate new ideas or varying outputs, as could be desired when generating images from prompts. During inference, we start with using randomly generated noise for our input and the condition, which is the RBG image taken from a different perspective, providing the information needed to replicate the pose seen in it. Predicting the desired result allows us to utilise only 50 denoising steps, significantly accelerating the inference. Our model leverages both the information provided by the condition and that acquired during later steps of the iterative denoising process to generate its output.
@@
Both paths follow the same structure of convolution layers, attention layers and maxpooling layers. The first path ends in two convolution layers that do not exist in the second downward path. The upward path consists of convolutional layers, self-attention layers and upsampling layers in order to get back to the dimensions of the original input. The latent connections from the second downward path to the upward path always go from the self-attention layer to the upsampling layers. We connect the encoding of the condition to the decoder with latent connections so that more spatial information from earlier layers is kept. The time step is encoded and given in the linear layers of the “Down” and “Up” modules.
@@
The network was trained for 100 epochs. As previously mentioned, the noise applied to each training sample was selected randomly from 1000 possible levels. This extends the variance of our dataset, because the label remains the same regardless of the level of noise applied to the input, and the model should consistently predict this label. The model was trained with the Adam optimiser.
@@
In this end-to-end approach, the network manages to identify the three-dimensional pose of the robot from a two-dimensional RBG image and to transpose that pose from a third-person perspective to an egocentric perspective, which is expressed through a generated RGB image.
III-B Datasets
Simulated NICOL. To create our dataset, we made use of the NICOL simulation introduced by Kerzel et al. [25]. We use two cameras in the simulation, one situated in the head of the robot and the second one positioned in front of it at a distance that allows us to fully capture the robot. The robot’s head is tilted downward to ensure the actions are captured in the first-person perspective. The NICOL robot is a semi-humanoid robot without legs integrated into a collaborative workspace. We make the robot randomly assume poses for both of its arms. In each of the kinematic chains representing the arms, we have 13 joints. Five of these joints are in the hands, while the other eight are in the arms. The joint values range from minus pi to pi, which we scale to be between 0 and 1 for the training. After the poses were assumed, one image from both cameras was captured and the joint values of the robot were recorded. The end-effector position and rotation are also recorded in seven values for each of the hands, with the first three values representing the x, y, and z coordinates and the other four values representing the orientation. During the data collection, 10000 samples were recorded. For training, 80% of the dataset was used, while 20% was used for validation. All the images are RBG and have a height and a width of 64 pixels. The random poses come from a uniform distribution but are limited to stay in the visual range and within the ranges of the joints. There are no images where the arms are behind the robot’s back or otherwise obscured. The finger joints are always kept in the same position, so only eight of the thirteen recorded joint values per arm are changed.
EMIL. The EMIL dataset [3] was collected with the NICO [26] robot. During the recording, NICO sat at a table and used its right arm to interact with an object on the table. There are four kinds of interactions: lifting the object, pushing the object, pulling the object or scooting the object to the side. Throughout these interactions, data from NICO and from external sensors was recorded. While data from multiple modalities is recorded, only the visual data from its camera and the external camera are relevant for our approach. The two visual streams are synchronised so that each frame recorded from NICO’s camera has a corresponding frame recorded from the external camera. In total, we used 59365 frames from each camera.
RH20T. The RH20T dataset [4] was collected with various robots positioned on a data collection platform. Multiple cameras, microphones, and other sensors collected data while the robots performed various tasks. The background is kept diverse through multiple table covers with complex colour patterns. The RH20T dataset contains demonstrations of more than 140 different tasks. We made use of the frames collected by two of the cameras (cam_036422060909, cam_104122063550) that record the robot from different perspectives. We used frames from the RH20T_cfg3 subset for a total of 187045 frames.